OpenStudy (anonymous):

Two boys want to use a seesaw, and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work. Can someone please help?

OpenStudy (confluxepic):

\[w_{1}\times d_{1}=w_{2} \times d_{2}\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

ok

OpenStudy (confluxepic):

\[d_{2}=\frac{ w_{1} \times d_{1} }{ w_{2} }\]

OpenStudy (confluxepic):

So that would be your final answer.

OpenStudy (anonymous):

thank you